OpenStudy (anonymous):

w^2+2w-8=0 solve by factoring

Directrix (directrix):

What are two numbers that multiply to -8 AND add to 2 ?

OpenStudy (anonymous):

-2 and 4

OpenStudy (anonymous):

Ao now substtute it dat way... Its a quadratic u no...

Directrix (directrix):

w^2+2w-8 = w^2 + 4w - 2w -8 = w(w + 4) - 2(w + 4) = (w - 2) * (w + 4). That is factoring by grouping.

Directrix (directrix):

(w - 2) * (w + 4) = 0 So, w-2 = 0 or w + 4 = 0 Solve for w @higherlearning

OpenStudy (anonymous):

2 and -4

Directrix (directrix):

Correct.